OpenStudy (thewalkingdeadrick):

Mr. Bernard owns a local pizza shop called Pizza Paradise that serves specialty pizzas in the restaurant and for delivery. Score Due to the popularity of the pineapple bacon pizza, Mr. Bernard has decided to make a larger size of the pizza. The diameter of the small pizza was 10 in. The diameter of the new, large pizza is 13 in. Anthony says the area of the larger pizza is 30% bigger because the diameter is 30% greater. Philip disagrees and says he thinks the area of the pizza is almost 70% bigger. Who is correct? Explain how you know and show your work. Use 3.14 for pi.

OpenStudy (thewalkingdeadrick):

will fan and medal

rebeccaxhawaii (rebeccaxhawaii):

calculate the area

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (mathstudent55):

Find the areas of the two pizzas. Then divide the larger area by the smaller area.

OpenStudy (thewalkingdeadrick):

so is it 40.82/31.4

OpenStudy (anonymous):

Correct:)

OpenStudy (thewalkingdeadrick):

thank you

OpenStudy (anonymous):

Can I have a medal not that I need one?

OpenStudy (anonymous):

Thank you!!!

OpenStudy (thewalkingdeadrick):

yw:)

OpenStudy (mathstudent55):

To find the areas you need the radius (half the diameter), and you need the formula for the area of a circle. 10-in. pizza d = 10 in.; r = 5 in. A = pi * r^2 = 3.14 * (5)^2 = 78.5 in.^2 13-in. pizza d = 13 in.; r = 6.5 in. A = pi * r^2 = 3.14 * (6.5)^2 = 132.7 in.^2 132.7/78.5 = 1.69 Since the large area divided by the small area is 1.69, the large pizza is 69% larger than the small pizza.